It costs 50,000 pesos at the end of each year to maintain a section of Kennon road in Baguio City. If money is worth 10%, how much would it pay to spend immediately to reduce the annual cost to 10,000 pesos?
Answer #1

Amount to pay to spend immediately to reduce the annual cost to 10,000 pesos is = p = A ÷ I

We're p = sum of money at present

A = series of periodic equal amount of payments

I = intrest rate

= (50,000 - 10,000) ÷ 10℅

= 40,000 × 100 / 10

= 4,00,000